OpenStudy (anonymous):

If log 2= a and log 3= b, express ln8 / ln5 in terms of a and b.

OpenStudy (anonymous):

ln 8 =ln 2^3-----)a^3 ln 5=ln3+ln2-----)a+b so,finally ...a^3/a+b

OpenStudy (anonymous):

i think \[\ln(5)\neq \ln(3)+\ln(2)\]

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

i think ln 3 + ln 2 should be ln 6... i actuallly got the answer ..it is a+2b / 1-a but i dont know the steps...

OpenStudy (anonymous):

ooops , sorry the question is ln 18 / ln 5

OpenStudy (anonymous):

Because the log are to base e they can be added like so ln2+ln3=ln(2*3)=ln6 Your numerator a+2b=ln18 is right but 1-a doesn't equal ln5